OpenStudy (anonymous):

Charlie was carrying around the steps to solve Square root of 6x plus 4 = 8. On his way through the Math Factory, Charlie dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 6x plus 4 = 8. A. Divide by 2 from both sides B. Square root both sides C. Divide by 6 from both sides D. Subtract 4 from both sides E. Subtract 8 from both sides F. Square both sides

OpenStudy (johnweldon1993):

\[\large \sqrt{6x + 4} = 8\] First would be to square both sides of this equation to eliminate the square root \[\large 6x + 4 = 64\] now we want to subtract 4 from both sides to get 6x by itself \[\large 6x = 60\] and last you should know...so what is the order?

OpenStudy (anonymous):

divide six from both sides leaving you with x=10

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

Thanks a bunch :D